Carl is 3 times as old as Sammy now. In 12 years, Carl will be twice as old as Sammy. Find their present
ages.

Answer by ikleyn(52788)

.

x for Sammy age


3x for Carl's age


In 12 years, as the problem states,


    2*(x+12) = 3x + 12


Simplify and solve


    2x + 24 = 3x + 12

    24 - 12 = 3x - 2x

       12   = x.


ANSWER.  Sammy is 12 years old now.  Carl is 3*12 = 36 years old.

Solved, answered and carefully explained.

----------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - HOW TO algebreze and to solve age problems?
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Age problem for three participants
    - Miscellaneous age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.